a)If the work function foe a metal is 1.85 ev, what would be the stopping potential for light having a wavelength of 410 nm? b) what would be the maximum speed of the emitted photoelectrons at the metals surface?


1
Expert's answer
2021-01-14T10:41:02-0500

By Enstein's photoelectric equation:


"KE_{max} = hc/\lambda-A"

where "KE_{max}" is the maximum kinetic energy of the photoelectrons, "h = 6.63\times 10^{-34}J\cdot s" is Plank's constant, "c = 3\times 10^{8}m/s" is the speed of light, "\lambda = 410nm = 4.1\times 10^{-7}m" is the wavelenght of the light, and "A = 1.85eV = 1.6\times 10^{-19}J" is the work function. On the other hand, the kinetic energy is:


"KE_{max} = eV_s"

where "e = 1.6\times 10^{-19}C" is the elementary charge, and "V_s" is the stopping potential. Thus, obtain:


"eV_s= hc/\lambda-A\\
V_s = \dfrac{ hc/\lambda-A}{e}\\
V_s = \dfrac{ 6.63\times 10^{-34}\cdot 3\times 10^8/(4.1\times 10^{-7})-1.6\times 10^{-19}}{1.6\times 10^{-19}}\\
V_s \approx 2.03V"

Since "KE_{max} = \dfrac{mv_{max}^2}{2}" (assuming classical limit), the maximum speed is:


"v_{max} = \sqrt{\dfrac{2\cdot KE_{max}}{m}} = \sqrt{\dfrac{2 eV_s}{m}}"

where "m = 9.1\times 10^{-31}kg" is the mass of the electron. Thus, obtain:


"v_{max} = \sqrt{\dfrac{2 \cdot 1.6\times 10^{-19}\cdot 2.03}{9.1\times 10^{-31}}} \approx 7.16\times 10^5m/s"

Answer. a) "2.03V" , b) "7.16\times 10^5m/s".